OpenStudy (anonymous):

distance between directrices equal to 9 square root 2 / 2, latus rectum equal to 2/3. find the equation of ellipse, center at origin.

OpenStudy (anonymous):

i am beginning to hate conic sections

OpenStudy (anonymous):

me too..

OpenStudy (anonymous):

but i have no choice..

OpenStudy (anonymous):

it has been months for this i think we can do this, but it will take me a minute or two

OpenStudy (anonymous):

this would be last topic before our midterm exam.

OpenStudy (anonymous):

okay thanks!! :)

OpenStudy (anonymous):

still here lets try this, hope i don't get lost

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

damn wrong already. i thought the directrix was given, but it is not back to the drawing board

OpenStudy (anonymous):

oh maybe it is one more piece of paper

OpenStudy (anonymous):

when a question includes distance between directrices i got stuck up..

OpenStudy (anonymous):

ok one more time directrix is \(x=\frac{9\sqrt{2}}{4}\)

OpenStudy (anonymous):

since your are given the distance, which is double that then this means \[\frac{a}{e}=\frac{9\sqrt{2}}{4}\] and so \[a=\frac{9\sqrt{2}}{4}e\]

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

also \[\frac{2b^2}{a}=\frac{2}{3}\] and so \[\frac{b^2}{a}=\frac{1}{3}\] making \[b^2=\frac{a}{3}\] and also \[b^2=a^2(1-e^2)\] together these give \[\frac{a}{3}=a^2(1-e^2)\] so \[\frac{1}{3}=a(1-e^2)\]

OpenStudy (anonymous):

then since \(a=\frac{9\sqrt2}{4}\) the last equation becomes \[\frac{1}{3}=\frac{9\sqrt2}{4}(1-e^2)\]

OpenStudy (anonymous):

i can get now the real value of e then substitute it to the equation to get a. yeheeey

OpenStudy (anonymous):

yeah but i am having a hell of a time solving it

OpenStudy (anonymous):

oh actually i guess you just need \(e^2\)

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

ok well i hope you are on the right path, because i gotta run good luck!

OpenStudy (anonymous):

THANK YOU SO MUCH!!